# Interactive Distance Field Mapping and Planning to Enable Human-Robot Collaboration

### v1

@@ IV Proposed Framework
For the next process of IDMP (yellow block), we build and maintain our main representation, the Fused Field, generated with all prior training points $\mathcal{P}_{\{0,...,i-1\}}$. We pass $\mathcal{P}_{\{0,...,i-1\}}$ to the Frustum Field to select points only within the frustum area, denoted as $\mathcal{P}_{f\{0,...,i-1\}}$. Then we use $\mathcal{P}_{f\{0,...,i-1\}}$ to query the Frustum Field for the distance $\hat{d}_{f}$ and gradient $\nabla\hat{d}_{f}$ as per Eq. (8). Based on the distance and gradient inference, the fusion method adjusts the training points’ positions, and the dynamic update method eliminates past training points of the objects moving in the scene. We call the intersection between $\mathcal{P}_{\{0,...,i-1\}}$ and $\mathcal{P}_{\{i\}}$ the overlapping area. Note that the new measurements that are out of the range of the overlapping area, denoted as $\mathcal{P}_{n\{i\}}$, are directly merged into the Fused Field. By updating the training points, which are maintained via an Octree-based data structure, we implicitly update the Fused Field.
For planning, let us define workspace points along the robot’s links as $\mathcal{P}_{o}$. The motion planning algorithm then queries the Fused Field for $\hat{d}_{o}$, $\nabla\hat{d}_{o}$ at the locations $\mathcal{P}_{o}$. These distances and gradients are used to compute a safe trajectory and can be repeatedly queried to adapt to a changing map.
@@ IV-B3 New Points Insertion
IV-C Motion Planning
Due to the differentiability of the Fused Field, our framework naturally accommodates gradient-based motion planners. For example, the motion planner can query $\hat{d}_{o}$ and $\nabla\hat{d}_{o}$ and utilise these to adapt its motion plans in response to a changing map. A key advantage of our framework is that we can use the gradient inference in Eq. 9 to compute the gradients analytically without surface normals, whereas conventional approaches rely on discrete grid-based approximation [8, 34]. This results in faster and more accurate computation of gradients which is crucial for close-proximity human-robot collaborative tasks where fast replanning is required.
@@ V-A Distance Field Mapping
The experiments are conducted on a laptop equipped with an AMD Ryzen 7 4800u CPU, 16GB RAM and no GPU. All three algorithms are allowed to multithread on up to 16 threads. Each algorithm is allowed to run as a ROS node with only the playback of the dataset (rosbag) running in the background. Original code implementations for Fiesta [^2] https://github.com/HKUST-Aerial-Robotics/FIESTA and Voxblox [^3] https://github.com/ethz-asl/voxblox were used for the experiments with default parameters.
Root Mean Squared Error (RMSE) is used to evaluate the distance field performance. For gradients, we use
@@ V-B1 Static Scene
(a) Ground truth
(b) IDMP output
(c) IDMP error
Figure 5: Evaluation of the IDMP distance field accuracy using a volume of size $4m$ x $2.75m$ x $1.25m$ from the Cow & Lady dataset. Queries are made at a regular grid every $5cm$.
Fig. 5 illustrates the true distance field, the IDMP output, and the absolute error of IDMP on three perpendicular planes on the Cow & Lady dataset. The planes are part of a volume that contains a mixture of complex surfaces but also free space with no obstacles.
Figure 6: Evaluation of distances and gradients using the Cow & Lady dataset. IDMP is compared to FIESTA and Voxblox. The evaluated volume is the same as in Fig. 5.
To compare IDMP’s distance field accuracy to Voxblox and FIESTA, we compute the RMSE for each framework while varying the spatial resolution. For Voxblox/FIESTA, this corresponds to the voxel size. For IDMP we query at a regular grid that corresponds to the centre of each voxel.
Figure LABEL:fig:cow_rmse shows our experimental results. The accuracy for Voxblox and FIESTA decreases with increasing spatial resolution. In contrast, IDMP’s accuracy remains constant and there is no requirement to pre-set a certain resolution illustrating the advantage of using a continuous representation.
The evaluation of gradients is presented in Fig. LABEL:fig:cow_cos. This figure shows the mean of the cosine similarities between the computed gradients and the ground truth. Both IDMP and FIESTA perform better than Voxblox for all voxel sizes and resolutions. FIESTA shows slightly better performance than IDMP in some cases; however, at the cost of higher variability. In terms of computation time for IDMP, updating the GP takes on average $250ms$ per frame for this scene while the query step takes $2\mu s$ per point.
@@ V-B2 Dynamic Scene
In this experiment, we evaluate how each framework handles dynamic scenes with moving objects. Fig. 6(a) shows the Gazebo scene with a ball rolling on the table from left to right. The last three subfigures show the surface points at the end of the run for the three frameworks. Both FIESTA and Voxblox show artefacts due to the progressive weighting and integration while IDMP does not. IDMP results in the best RMSE of $2.6cm$ which is more than a factor of $2$ better than FIESTA/Voxblox. Updating the GP for a scene of this size takes on average $50ms$ per frame while the query step takes $2\mu s$ per point. Only the time for processing the frame is dependent on the size of the scene whereas the time needed for the query remains constant.
@@
Figure 7: Comparison of IDMP with FIESTA and Voxblox for a dynamic scene of size $0.39m$ x $1.05m$ x $0.35m$. (a) shows the setup in simulation while (b) to (d) show the fused surface points at the end of the run. Colours indicate the height of the points. The spatial resolution is $1cm$. IDMP handles moving objects seamlessly while FIESTA and Voxblox keep artefacts from older frames.
V-C Motion Planning in Dynamic Scenes
In this experiment we demonstrate that our method is able to facilitate safe motion planning of a robot arm in the presence of dynamic obstacles which is useful for close-proximity human-robot collaboration tasks. To simulate such a scenario we task the robot with moving to a goal pose while avoiding a moving ball object, see Fig. 8.
Trajectory optimisers are well suited to such scenarios due to their fast computation and smooth trajectory properties. For our experiment we utilise the Covariant Hamiltonian optimisation-based motion planner (CHOMP) [8], a gradient-based trajectory optimisation method that iteratively perturbs a given initial trajectory away from obstacles. To do so, it utilises the gradient field to minimise its optimisation objective which consists of trajectory smoothness and obstacle avoidance terms. Due to a) the ability of our framework to compute updated gradients analytically and b) the need to only query the points $\mathcal{P}_{o}$ the motion planner requires, we can continuously replan online. The computation time for the query shown in Fig. 8 is approximately $6.4ms$ ($3190$ points at $2\mu s/$point).
@@ V-C Motion Planning in Dynamic Scenes
Figure 8: Re-planning in the presence of a dynamic obstacle. CHOMP approximates the manipulator via $29$ spheres. For the trajectory shown, $110$ points per sphere are used, resulting in $29*110=3190$ query points.

### v3

@@ IV Proposed Framework
The next process of IDMP (yellow block) builds and maintains our main representation, the Fused Field, generated with all prior training points $\mathcal{P}_{\{0,...,i-1\}}$. We pass $\mathcal{P}_{\{0,...,i-1\}}$ to the Frustum Field to select points only within the frustum area, denoted as $\mathcal{P}_{f\{0,...,i-1\}}$. Then we use $\mathcal{P}_{f\{0,...,i-1\}}$ to query the Frustum Field for the distance $\hat{d}_{f}$ and gradient $\nabla\hat{d}_{f}$ as per Eq. (8). Based on the distance and gradient inference, the fusion method adjusts the training points’ positions, and the dynamic update method eliminates past training points of the objects moving in the scene. We call the intersection between $\mathcal{P}_{\{0,...,i-1\}}$ and $\mathcal{P}_{\{i\}}$ the overlapping area. Note that new measurements that are out of the range of the overlapping area, denoted as $\mathcal{P}_{n\{i\}}$, are directly merged into the Fused Field. By updating the training points, which are maintained via an Octree-based data structure, we update the Fused Field.
For planning, let us define workspace points along the robot’s links as $\mathcal{P}_{o,s}$. The motion planning algorithm then queries the Fused Field for $\hat{d}_{o,s}$, $\nabla\hat{d}_{o,s}$ at $\mathcal{P}_{o,s}$. These distances and gradients are used to compute a safe trajectory and can be repeatedly queried to adapt to a changing map.
@@ IV-B3 New Points Insertion
IV-C Reactive Planning
Due to the differentiability and efficiency of the Fused Field, our framework naturally accommodates reactive planning. For example, a gradient-based motion planner can query $\hat{d}_{o}$ and $\nabla\hat{d}_{o}$ and utilise these to adapt its motion plans in response to a changing map. A key advantage of our framework is that we can use the gradient inference in Eq. 9 to compute the gradients analytically, whereas conventional approaches rely on discrete grid-based approximation [7]. This results in faster and more accurate computation of gradients which is crucial for close-proximity human-robot collaborative tasks where fast replanning is required.
In addition, our framework naturally accommodates fast local reactive methods. Consider the current point $\mathbf{x}_{s}$ and a goal point $\mathbf{x}_{g}$. Given the queried distance and gradient of the current point as $\hat{d}_{s}$ and $\nabla\hat{d}_{s}$ from our IDMP Fused Field, we can formulate a repulsive vector $\mathbf{v}_{rep}=\nabla\hat{d}_{s}$ and an attractive vector $\mathbf{v}_{att}=\mathbf{x}_{g}-\mathbf{x}_{s}$. Note that both vectors need to be normalised. The repulsive behaviour is defined by a function $w(\hat{d}_{s})$ dependent on the distance. We then can formulate the resulting vector by combining the repulsive vector and attractive vector using $w(\hat{d}_{s})$ as in [33],
$\mathbf{v}_{res}=w(\hat{d}_{s})\mathbf{v}_{rep}+(1-w(\hat{d}_{s}))\mathbf{v}_{att}.$ (11)
By efficiently computing Eq. 11 at each timestep, a robot has the ability to follow the resulting vector to reactively avoid moving objects present in the scene in an online manner.
@@ V-A Distance Field Mapping
The experiments are conducted on a laptop equipped with an AMD Ryzen 7 4800u CPU, 16GB RAM and no GPU. All three algorithms are allowed to multithread on up to 16 threads. Each algorithm is allowed to run as a ROS node with only the playback of the dataset (rosbag) running in the background. Original code implementations for Fiesta [^2] https://github.com/HKUST-Aerial-Robotics/FIESTA and Voxblox [^3] https://github.com/ethz-asl/voxblox were used for the experiments with default parameters. Root Mean Squared Error (RMSE) is used to evaluate the distance field performance. For gradients, we use
@@ V-B1 Static Scene
Figure 5: Quantitative evaluation of distance RMSE (lower is better), gradients cosine similarity (higher is better) and computation time (lower is better) using the Cow & Lady dataset for IDMP, FIESTA and Voxblox.
To compare IDMP to Voxblox and FIESTA, we compute the RMSE for each framework while varying the spatial resolution. For Voxblox/FIESTA, this corresponds to the voxel size. For IDMP we vary the resolution of the training points used to update the Fused Field. Figure 5 shows our experimental results. It shows the distance and gradient accuracy, and the overall time needed to process a single frame.
The accuracy for Voxblox and FIESTA decreases with increasing spatial resolution. In contrast, IDMP’s accuracy remains constant, outperforming FIESTA with resolutions higher than 15cm. Note that even though we increase the resolution of our training points, IDMP generates a continuous distance field, that can be queried at any point in space independent of the spatial resolution. Voxblox and FIESTA only generate a discrete distance field in the predefined spatial resolution and where data was observed.
The evaluation of gradients shows the mean of the cosine similarities between the computed gradients and the ground truth. Both IDMP and FIESTA perform better than Voxblox for all spatial resolutions. FIESTA shows slightly better performance than IDMP in some cases; however, at the cost of higher variability.
The last part shows the average computation time for each algorithm to process a frame. For IDMP this includes generating the Frustum Field, performing the fusion, updating dynamic objects and generating the Fused Field. For Fiesta and Voxblox this includes the ESDF/TSDF generation, integration and raycasting. IDMP outperforms Voxblox for every spatial resolution while reaching similar computation times as FIESTA for resolutions larger than $15$ cm.
Regarding query times, it should be noted that while Voxblox and Fiesta have constant distance and gradient query times, IDMP has a small overhead time ($2\mu$s per point). However, the benefit of our method is that queries can be made at any arbitrary resolution post-mapping due to our continuous representation.
@@ V-B2 Dynamic Scene
In this experiment, we evaluate how each framework handles dynamic scenes with moving objects. Fig. 6(a) shows the Gazebo scene with a ball rolling on the table from left to right. The last three subfigures show the surface points at the end of the run for the three frameworks. Both FIESTA and Voxblox show artefacts due to the progressive weighting and integration, while IDMP does not. IDMP results in the best RMSE of $2.6$ cm which is more than a factor of $2$ better than FIESTA/Voxblox. Updating the GP for a scene of this size takes on average $50$ ms per frame while the query step takes $2~{}\mu$s per point. Only the time for processing the frame is dependent on the size of the scene, whereas the time needed for the query remains constant.
@@
Figure 6: Comparison of IDMP, FIESTA and Voxblox for a dynamic scene. (a) Shows the stimulation setup, and (b) to (d) show the fused surface points at the end of the sequence. Colours indicate the point’s height with a spatial resolution of 1 cm. IDMP handles appropriately moving objects while FIESTA and Voxblox keep artefacts from older frames.
V-C Reactive Planning in Dynamic Scenes
In this experiment, we demonstrate that our method is able to facilitate safe reactive motion and path planning of a robot arm in the presence of dynamic obstacles, which is useful for close-proximity human-robot collaboration tasks. To simulate such a scenario, we task the robot with moving to a goal pose while avoiding a moving ball object, see Fig. 7.
Trajectory optimisers are well suited to such scenarios due to their fast computation and smooth trajectory properties. We utilise CHOMP [7], a gradient-based trajectory optimisation method that iteratively perturbs a given initial trajectory away from obstacles. To do so, it utilises the gradient field to minimise its optimisation objective, which consists of trajectory smoothness and obstacle avoidance terms. Due to a) the ability of our framework to compute updated gradients analytically and b) the need to only query the points $\mathcal{P}_{o}$ the motion planner requires, we can continuously replan online. The computation time for the query shown in Fig. 7 is approximately $6.4$ ms ($3190$ points at $2~{}\mu$s point).
We additionally validate IDMP in combination with CHOMP in a real-world scenario where a human moves into the robot’s workspace mid-way during its trajectory (see Fig. 8). Furthermore, we demonstrate the proposed local reactive method in a similar scenario. Fig. 9 shows the repulsive vector in red, the attractive vector in blue and the resulting vector in green. The robot arm follows the resulting vector to quickly and safely avoid the human’s moving arm. Because the repulsive vector only uses the information at the robots current position, querying the Fused Field takes $2~{}\mu$s, enabling real-time obstacle avoidance.
@@ V-C Reactive Planning in Dynamic Scenes
Figure 7: Re-planning in the presence of a dynamic obstacle. CHOMP approximates the manipulator via $29$ spheres. For the trajectory shown, $110$ points per sphere are used, resulting in $29\times 110=3190$ query points.
Figure 8: Real-world experiment of IDMP with CHOMP. Re-planning in the presence of a human arm.
Figure 9: IDMP and local reactive obstacle avoidance.